OpenStudy (anonymous):

If 5p = -100 + 3Q2, which of the following could Q possibly be?

OpenStudy (anonymous):

A. Q = square root of 5 times p minus 100 over 3 B. Q = square root of 5 times p plus 100 over 3 C. Q = square root of negative 5 times p plus 100 over 3 D. Q = square root of negative 5 times p minus 100 over 3

OpenStudy (mathmale):

What do you mean by "Q2?"

OpenStudy (anonymous):

5p = -100 +3Q^2 sorry

OpenStudy (mathmale):

So much clearer. Thank you. It appears your job is to solve the given equation for Q. How would you do that?

OpenStudy (anonymous):

I'm not really sure, I think I should subtract -100 from 5p

OpenStudy (anonymous):

Which would really just be adding it

OpenStudy (mathmale):

Why not "add 100 to both sides?"

OpenStudy (anonymous):

Yeah that's what I meant haha

OpenStudy (anonymous):

Then it would be 5p +100 = 3Q^2 so then I would divide both sides by 3

OpenStudy (mathmale):

That's right. And what would your final solution step be?

OpenStudy (anonymous):

(5p+100)/3 = Q^2

OpenStudy (anonymous):

So square root of (5p+100)/3 = Q

OpenStudy (mathmale):

Good use of parentheses! How many answers would you then have?

OpenStudy (anonymous):

What do you mean?

OpenStudy (mathmale):

If Q^2=32, then you have two possible answers, one positive and one negative:\[Q=+\sqrt{32}~U~Q=-\sqrt{32}\] or just "plus or minus Sqrt(32)."

OpenStudy (mathmale):

Two answers, correct?

OpenStudy (anonymous):

I don't understand how you got that

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

But wouldn't my answer for this problem just be B. Q = square root of 5 times p plus 100 over 3